OpenStudy (anonymous):

A box has 5 beads of the same size, but all are different colors. Tina draws a bead randomly from the box, notes its color, and then puts the bead back in the box. She repeats this 3 times. What is the probability that Tina would pick a red bead on the first draw, then a green bead, and finally a red bead again? A. 1/625 B. 1/180 C. 1/150 D. 1/125

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

Since the bead is replaced after each draw, they are independent events. As in the last question, the overall probability is the product of the individual probabilities for each draw, i.e.\[P(RGR) = P(R)\times P(G)\times P(R)\]Because the 5 beads are all different colors, you can easily see what the probability of drawing each color on a single draw is.

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

so D is right? @ospreytriple ?

OpenStudy (anonymous):

since 5*5*5 is 125

OpenStudy (anonymous):

Correct. Well done.